# Matryoshka: Learning to Drive Black-Box LLMs with LLMs

### v1

@@ Abstract
Despite the impressive generative abilities of black-box large language models (LLMs), their inherent opacity hinders further advancements in capabilities such as reasoning, planning, and personalization. Existing works aim to enhance LLM capabilities via domain-specific adaptation or in-context learning, which require additional training on accessible model parameters, an infeasible option for black-box LLMs. To address this challenge, we introduce Matryoshka, a lightweight white-box LLM controller that guides a large-scale black-box LLM generator by decomposing complex tasks into a series of intermediate outputs. Specifically, we consider the black-box LLM as an environment, with Matryoshka serving as a policy to provide intermediate guidance through prompts for driving the black-box LLM. Matryoshka is trained to pivot the outputs of the black-box LLM aligning with preferences during iterative interaction, which enables controllable multi-turn generation and self-improvement in optimizing intermediate guidance. Empirical evaluations on three diverse tasks demonstrate that Matryoshka effectively enhances the capabilities of black-box LLMs in complex, long-horizon tasks, including reasoning, planning, and personalization. By leveraging this pioneering controller-generator framework to mitigate dependence on model parameters, Matryoshka provides a transparent and practical solution for improving black-box LLMs through controllable multi-turn generation using white-box LLMs.
@@ 1 Introduction
Most of the commercial large language models (LLMs) (Radford et al., 2019; Brown, 2020; Achiam et al., 2023; Chowdhery et al., 2023; Team et al., 2023; Reid et al., 2024) are black-box models (Sun et al., 2024b; Zhuang et al., 2024b), where the model structure, parameters, or even output logits are not accessible. Although these black-box LLMs have exhibited remarkable efficacy across a diverse array of applications, revolutionizing natural language processing tasks such as text completion (Radford et al., 2019; Brown, 2020), translation (Zhu et al., 2023), question-answering (Hendrycks et al., 2020), etc, the applications of black-box LLMs continue to face significant challenges when faced with tasks that require more advanced cognitive capabilities, particularly in the realms of reasoning (Hendrycks et al., 2021; Wang et al., 2024b), planning (Valmeekam et al., 2022; Zhuang et al., 2023; Jimenez et al., 2023; Mialon et al., 2023; Zhuang et al., 2024a; Shi et al., 2024b), and personalization problems (Salemi et al., 2023; Tan et al., 2024a). Enhancing such capabilities within black-box LLMs presents unique challenges, primarily due to the lack of direct access to internal model parameters (Huang et al., 2023; Sun et al., 2024b; Zhuang et al., 2024b). This opacity introduces substantial complexity in efforts to refine and augment these advanced cognitive functions within the framework of black-box architectures.
@@
Existing research efforts for improving black-box LLM performance can be largely categorized into two main methodological paradigms (Figure 1): (1) In-context learning (ICL)-based methods (Sun et al., 2024a; Tan et al., 2024b; Zhuang et al., 2024b) that are designed to guide LLM in exhibiting specific capabilities or adhering to particular directives. However, these frameworks necessitate meticulously constructing few-shot demonstrations or prompts for LLMs to emulate or follow, rather than fundamentally advancing their intrinsic capabilities. (2) Adapter-based methods (Sun et al., 2024b; Zhuang et al., 2024b; Shi et al., 2024a) that exploit the inherent randomness in LLM generation, producing multiple candidate outputs and subsequently selecting those that optimally satisfy domain-predetermined criteria. Nevertheless, these approaches are highly dependent on the intrinsic synthetic capabilities or built-in functionalities of the black-box LLM, potentially resulting in the selection of a suboptimal candidate when all the generated options are less than ideal. Furthermore, both ICL and adapter-based methodologies exhibit significant limitations when applied to long-horizon tasks (e.g., multi-step reasoning, long-term planning, etc.) due to their inherent lack of environmental interaction capabilities. In light of these constraints, we propose to leverage smaller, open-source LLMs as controllers to generate soft prompts as guidance, instead of relying on hard memory in context.
@@
Figure 1: Enhancement in black-box LLMs capabilities. Existing methods either (a) integrate well-crafted instructions or meticulously-picked few-shot demonstrations as guidance or (b) exploit randomness in model generations to identify the most promising solution from candidates. In Matryoshka, we present (c) a controller-generator framework that enables white-box LLMs to drive the behavior of black-box LLMs for enhanced capabilities. Refer to caption indicates the trainable parameters, whereas Refer to caption indicates the inaccessible fixed parameters.
@@
Similar to the scratchpad in o1-preview (OpenAI, 2024c) [^1] Scratchpad is a sequence of intermediate chain-of-thoughts generated prior to producing the final answer. In Matryoshka, we broaden the definition of intermediate tokens to encompass various forms of guidance that can enhance the capabilities of LLMs, including task decomposition and user history summarization., we propose Matryoshka, a modular framework designed to enhance the advanced problem-solving capabilities of black-box LLMs via controllable multi-turn generations. Matryoshka consists of a lightweight white-box LLM that functions as a controller and a black-box LLM that serves as a generator or solver. Upon receiving the question description as input, the controller generates intermediate outputs that augment the capabilities of the subsequent black-box LLMs. For example, the controller can decompose the original complex task into high-level subtasks in reasoning or planning scenarios, or summarize profiles from historical records for personalization tasks. By conceptualizing the following black-box LLM as the environment, Matryoshka generates intermediate guidance alongside the original input to derive the final result through multi-turn interactions with the environment. The feedback for the outputs from the environments distinguishes positive and negative examples of intermediate generations, which can be used for preference optimization. Notably, this optimization process is inherently self-improving through iterative sampling from prior inferences and by considering the policies from earlier iterations as reference policies. Matryoshka continually enhances the advanced capabilities of the black-box LLM through controllable multi-turn generations that iteratively interact with environmental feedback.
@@
Extensive experiments conducted on three complex tasks demonstrate the effectiveness and generalizability of Matryoshka in improving the advanced problem-solving capabilities of black-box LLMs, with an average improvement of 3.19% in accuracy for reasoning, 7.46% in success rate for planning, and 5.82% in accuracy for personalization. Importantly, Matryoshka not only enhances the capabilities of black-box LLMs without requiring access to model parameters, but also facilitates online feedback with environmental interactions. We summarize the main contributions as follows:
@@
  • i), We introduce Matryoshka, one of the first modular frameworks that employ a lightweight white-box LLM to drive the generation of a large-scale black-box LLM for complex problem-solving;
@@
  • ii), Matryoshka intuitively formulates the white-box LLM as a controller and the black-box LLM as a component of the environment, facilitating long-horizon controllable generation with environmental feedback; and
@@
  • iii), Matryoshka adopts on-policy learning to iteratively enhance training data quality, inherently self-improving intermediate guidance for the continual enhancement of black-box LLM capabilities.
@@ 2 Problem Formulation
Our objective is to enhance the capability of a black-box LLM in solving complex, long-horizon problems by calibrating its output generation to better align with specific tasks. To achieve this, we conceptualize both the original outputs and the optimal solutions as distributions within a joint space, $\mathcal{Y}\sim\mathcal{Y}^{\text{org}}\times\mathcal{Y}^{\text{sol}}$, where $\mathcal{Y}^{\text{org}}$ and $\mathcal{Y}^{\text{sol}}$ represent the original text generations and target solutions, respectively. Specifically, given a set of task descriptions $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$, our goal is to adjust the outputs $\hat{y}_{i}\in\mathcal{Y}^{\text{org}}$ of the black-box LLM toward the hidden target solutions $y_{i}\in\mathcal{Y}^{\text{sol}}$ that successfully solve the problems. This involves driving the black-box LLM to generate outputs more closely aligned with the desired solutions without requiring access to its internal parameters.
@@
Figure 2: Controller-generator framework in Matryoshka comprising a white-box LLM as the controller and a black-box LLM as the generator and part of the environment. Given an input query $x$, Matryoshka leverages the intermediate generation $f_{\theta}(x)$ from the controller $\theta$ to drive the generator’s behavior. The final answer is derived from the generation $y\sim g(f_{\theta}(x))$.
@@
White-Box LLM as Controller/Policy. We propose utilizing a lightweight white-box language model as a controller to enhance the capabilities of black-box LLMs in solving various tasks. The process begins by feeding a text-grounded task description $x$ from the task space $\mathcal{X}$ into a smaller language model $\theta$, which acts as the controller. This smaller model generates $f_{\theta}(x)$, an automatic prompt designed to augment the performance of black-box LLMs on the specific task. These prompts can facilitate various functions, such as chain-of-thoughts for reasoning, task decomposition for planning, and user profile summarization from historical records for personalization. The generated intermediate prompt $f_{\theta}(x)$ is then combined with the original problem description $x$ and input $(x,f_{\theta}(x))$ into the black-box LLM. The capability enhancement will be measured by evaluating the performance improvements achieved through this controller-generator framework.
We emphasize that unlike works focusing on token-level LLM policy (Wang et al., 2024a; Rafailov et al., 2024a), our action space consists of entire intermediate guidance generations to solve the task.
Black-Box LLM as Environment. We recognize the black-box LLM as an environment to be controlled by the white-box LLM policy. After inputing the prompts $(x,f_{\theta}(x))$, the black-box LLM produces a final solution $\hat{y}=g_{\text{LLM}}(x,f_{\theta}(x))$ for the task. We utilize the final correctness of the black-box LLM’s output to evaluate the quality $u(x,f_{\theta}(x))$ as the reward of the intermediate guidance produced by the white-box LLM controller (Figure 2):
$u(x,f_{\theta}(x)):=\text{eval}(x,g_{\text{LLM}}(x,f_{\theta}(x)),$ (1)
where $\text{eval}(\cdot)$ denotes the oracle evaluation function of the final answer. For example, in question-answering tasks with ground-truth final answer $y$, the evaluation function measures accuracy by comparing the prediction with ground truth as $\text{eval}(x,g_{\text{LLM}}(x,f_{\theta}(x))=\mathbbm{1}(g_{\text{LLM}}(x,f_{\theta}(x))=y)$, where $\mathbbm{1}(\cdot)$ is the indicator function. For planning tasks without a ground-truth solution, the evaluation function assesses the success rate after executing the final solution as $\text{eval}(x,g_{\text{LLM}}(x,f_{\theta}(x))=\mathbbm{1}_{\text{succ}}(g_{\text{LLM}}(x,f_{\theta}(x)))$.
@@
For initialization, a prompt $x$ is sampled from task space $\mathcal{X}$ and serves as the initial state $s_{0}=x$. At each subsequent step $t\in[T]$, the controller generates prompts $a_{t}$ based on the current $s_{t-1}$. In response to the controller’s action, the environment returns an observation $o_{t}$ based on the history $s_{t-1}$ and the current action $a_{t}$. The state then transitions to include the new action and observation:
@@
The framework formulates a Markov Decision Process (MDP), which offers the potential for solving tasks that require long-horizon generations, including long-term planning and multi-step reasoning. By obtaining feedback from $\text{eval}(\cdot)$, we can conduct multi-turn optimization over the white-box LLM controller on the intermediate generations. Additionally, the multi-turn interaction with the environment during the data sampling stage can help improve data quality. Although optimizing this guidance presents challenges due to the inaccessibility of the black-box LLM’s parameters that preclude backpropagation of gradients during training, the existing reinforcement learning techniques, e.g.Schulman et al. (2017); Rosset et al. (2024), can be used for policy optimization.
3 Matryoshka
In this section, we specialize the white-box LLM controller that generates intermediate guidance to assist in task understanding and problem-solving in Section 3.1 and discuss the data collection procedure by interacting with black-box LLM in Section 3.2, which will be used for Matryoshka training to align the outputs of the black-box LLM with preferences in Section 3.3.
@@ 3.1 Instantiation of White-Box LLM Controller
Figure 3: Examples of intermediate guidance generated by Matryoshka for complex reasoning, planning, and personalization tasks.
We instantiate the white-box LLM as a controller to generate additional guidance that assists the black-box LLM in understanding and solving a diverse range of problems. Given the varying complexity and distinct characteristics of different tasks, the controller should be capable of generating guidance in various formats. We provide examples corresponding to reasoning, planning, and personalization tasks (Figure 3):
Problem Decomposition for Reasoning. For reasoning tasks, generating a sequence of reasoning steps is essential to solve the problem effectively. Existing works (Zhou et al., 2023) have observed that models often perform poorly on tasks that require solving problems more complex than the exemplars provided in the prompts. To enable the model to develop better reasoning and overcome the easy-to-hard generalization issue, one strategy is to decompose complex problems into a series of simpler sub-problems and solve them sequentially. Therefore, for reasoning tasks, the white-box LLM controller outputs decomposed sub-tasks to assist the subsequent black-box LLM generator in enhancing its reasoning capabilities.
High-Level Plan for Planning. For planning tasks, LLMs are required to generate a sequence of actions that constitute a plan to solve the given problems. A common strategy (Sun et al., 2024a; Zhao et al., 2024) is to apply hierarchical planning for complex solutions, where a high-level planner decomposes the task into sub-goals, and a low-level planner generates a sequence of admissible actions corresponding to each specific sub-goal. To enhance the black-box LLM’s planning capabilities, we leverage the white-box controller to generate high-level plans as guidance to simplify the problems.
User History Summarization for Personalization. For personalization tasks, LLMs are required to tailor outputs to individual users. Existing work (Richardson et al., 2023) accomplishes this by concatenating the user’s input query with a profile summarizing the user’s preferences and behavior patterns. To enhance the black-box LLM’s personalization capabilities, we utilize the white-box LLM controller to generate summaries of user histories. This approach enables black-box LLMs to better understand users and generate tailored content accordingly.

### v2

# Matryoshka Pilot: Learning to Drive Black-Box LLMs with LLMs

@@ Abstract
Despite the impressive generative abilities of black-box large language models (LLMs), their inherent opacity hinders further advancements in capabilities such as reasoning, planning, and personalization. Existing works aim to enhance LLM capabilities via domain-specific adaptation, which require additional training on accessible model parameters, an infeasible option for black-box LLMs. To address this challenge, we introduce Matryoshka Pilot (M-Pilot), a lightweight white-box LLM controller that guides a large-scale black-box LLM generator by decomposing complex tasks into a series of intermediate outputs. Specifically, we consider the black-box LLM as an environment, with M-Pilot serving as a policy to provide intermediate guidance through prompts for driving the black-box LLM. M-Pilot is trained to pivot the outputs of the black-box LLM aligning with preferences during iterative interaction, which enables controllable multi-turn generation and self-improvement in optimizing intermediate guidance. Empirical evaluations on diverse tasks demonstrate that our method effectively enhances the capabilities of black-box LLMs in complex, long-horizon tasks.
@@ 1 Introduction
Most of the commercial large language models (LLMs) (radford2019language, ; brown2020language, ; achiam2023gpt, ; chowdhery2023palm, ; team2023gemini, ; reid2024gemini, ) are black-box models (sunbbox, ; zhuang2024hydra, ), where the model structure, parameters, or even output logits are not accessible. Although these black-box LLMs have exhibited remarkable efficacy across a diverse array of applications, revolutionizing natural language processing tasks such as text completion (radford2019language, ; brown2020language, ), translation (zhu2023multilingual, ), question-answering (hendrycks2020measuring, ), etc, the applications of black-box LLMs continue to face significant challenges when faced with tasks that require more advanced cognitive capabilities, particularly in the realms of reasoning (hendrycks2021measuring, ; wang2024mmlu, ), planning (valmeekam2022large, ; zhuang2023toolqa, ; jimenez2023swe, ; mialon2023gaia, ), and personalization problems (salemi2023lamp, ; tan2024democratizing, ). Enhancing such capabilities within black-box LLMs presents unique challenges, primarily due to the lack of direct access to internal model parameters (huang2023k, ; sunbbox, ; zhuang2024hydra, ). This opacity introduces substantial complexity in efforts to refine and augment these advanced cognitive functions within the framework of black-box architectures.
@@
Existing research efforts for improving black-box LLM performance can be largely categorized into two main methodological paradigms (Figure 1): (1) In-context learning (ICL)-based methods (sun2024adaplanner, ; tan2024democratizinglargelanguagemodels, ; zhuang2024hydra, ) that are designed to guide LLM in exhibiting specific capabilities or adhering to particular directives. However, these frameworks necessitate meticulously constructing few-shot demonstrations or prompts for LLMs to emulate or follow, which relies on heuristic prompts constructions. (2) Adapter-based methods (sunbbox, ; zhuang2024hydra, ; shi2024medadapter, ) that exploit the inherent randomness in LLM generation, producing multiple candidate outputs and subsequently selecting those that optimally satisfy domain-predetermined criteria. Nevertheless, these approaches are highly dependent on the intrinsic synthetic capabilities or built-in functionalities of the black-box LLM, potentially resulting in the selection of a suboptimal candidate when all the generated options are less than ideal. Furthermore, both ICL and adapter-based methodologies exhibit significant limitations when applied to long-horizon tasks (e.g., multi-step reasoning, long-term planning, etc.) due to their inherent lack of environmental interaction capabilities. In light of these constraints, we propose to leverage smaller, open-source LLMs as controllers to generate soft prompts as guidance, instead of relying on hard memory in context.
@@
Figure 1: Enhancement in black-box LLMs capabilities. Existing methods either (a) integrate well-crafted instructions or meticulously-picked few-shot demonstrations as guidance or (b) exploit randomness in model generations to identify the most promising solution from candidates. In M-Pilot, we present (c) a controller-generator framework that enables white-box LLMs to drive the behavior of black-box LLMs for enhanced capabilities. Refer to caption indicates the trainable parameters, whereas Refer to caption indicates the inaccessible fixed parameters.
@@
Similar to the scratchpad in o1-preview [^1] Scratchpad is a sequence of intermediate chain-of-thoughts generated prior to producing the final answer. In M-Pilot, we broaden the definition of intermediate tokens to encompass various forms of guidance that can enhance the capabilities of LLMs, including task decomposition and user history summarization. (o1-preview, ), we propose Matryoshka Pilot (M-Pilot), a modular framework designed to enhance the advanced problem-solving capabilities of black-box LLMs via controllable multi-turn generations. M-Pilot consists of a lightweight white-box LLM that functions as a controller and a black-box LLM that serves as a generator or solver. Upon receiving the question description as input, the controller generates intermediate outputs that augment the capabilities of the subsequent black-box LLMs. For example, the controller can decompose the original complex task into high-level subtasks in reasoning or planning scenarios, or summarize profiles from historical records for personalization tasks. By conceptualizing the following black-box LLM as the environment, M-Pilot generates intermediate guidance alongside the original input to derive the final result through multi-turn interactions with the environment. The feedback for the outputs from the environments distinguishes positive and negative examples of intermediate generations, which can be used for preference optimization. Notably, this optimization process is inherently self-improving through iterative sampling from prior inferences and by considering the policies from earlier iterations as reference policies. M-Pilot continually enhances the advanced capabilities of the black-box LLM through controllable multi-turn generations that iteratively interact with environmental feedback.
@@
Extensive experiments conducted on three complex tasks demonstrate the effectiveness and generalizability of M-Pilot in improving the advanced problem-solving capabilities of black-box LLMs, with an average improvement of 3.19% in accuracy for reasoning, 7.46% in success rate for planning, and 5.82% in accuracy for personalization. Importantly, M-Pilot not only enhances the capabilities of black-box LLMs without requiring access to model parameters, but also facilitates online feedback with environmental interactions. We summarize the main contributions:
@@
  • i), We introduce M-Pilot, one of the first modular frameworks that employ a lightweight white-box LLM to drive the generation of a large-scale black-box LLM for complex problem-solving;
@@
  • ii), M-Pilot intuitively formulates the white-box LLM as a controller and the black-box LLM as a component of the environment, facilitating long-horizon controllable generation with feedback;
@@
  • iii), M-Pilot adopts on-policy learning to iteratively enhance training data quality, inherently self-improving intermediate guidance for the continual enhancement of black-box LLM capabilities.
@@ 2 Problem Formulation
Our objective is to enhance the capability of a black-box LLM in solving complex, long-horizon problems by calibrating its output generation to better align with specific tasks. To achieve this, we conceptualize both the original outputs and the optimal solutions as distributions within a joint space, $\mathcal{Y}\sim\mathcal{Y}^{\text{org}}\times\mathcal{Y}^{\text{sol}}$, where $\mathcal{Y}^{\text{org}}$ and $\mathcal{Y}^{\text{sol}}$ represent the original text generations and target solutions, respectively. Specifically, given a set of task descriptions $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$, our goal is to adjust the outputs $\hat{y}_{i}\in\mathcal{Y}^{\text{org}}$ of the black-box LLM toward the hidden target solutions $y_{i}\in\mathcal{Y}^{\text{sol}}$ that successfully solve the problems. This involves driving the black-box LLM to generate outputs more closely aligned with the desired solutions without access to parameters.
@@
Figure 2: Controller-generator framework in M-Pilot comprising a white-box LLM as the controller and a black-box LLM as the generator and part of the environment. Given an input query $x$, M-Pilot leverages the intermediate generation $f_{\theta}(x)$ from the controller $\theta$ to drive the generator’s behavior. The final answer is derived from the generation $y\sim g_{\text{LLM}}(f_{\theta}(x))$.
@@
White-Box LLM Drives Black-Box LLMs. To enhance the capabilities of black-box LLMs in solving various tasks, we introduce a lightweight white-box language model as a controller. The process begins by feeding a text-grounded task description $x$ from the task space $\mathcal{X}$ into a smaller language model $\theta$, which acts as the controller. The smaller model generates a sequence of $T$-step intermediate guidance $\{g_{t}\}_{t=1}^{T}\sim f_{\theta}(x)$ to augment the performance of black-box LLMs on the specific task. These guidances can facilitate various functions, such as chain-of-thoughts for reasoning, task decomposition for planning, and user profile summarization from historical records for personalization. The generated intermediate guidance $\{g_{t}\}_{t=1}^{T}$ is then combined with the original problem description $x$ and input $(x,\{g_{t}\}_{t=1}^{T})$ into the black-box LLM to obtain the final prediction $\hat{y}\sim g_{\text{LLM}}(x,\{g_{t}\}_{t=1}^{T})$. To formally characterize this process, we define a trajectory as:
[…]
which encapsulates the task description $x$, the intermediate guidance sequence $\{g_{t}\}_{t=1}^{T}$, and the final prediction $\hat{y}$. Assuming an autoregressive generation process for both the intermediate guidance and final output, the conditional probability of the trajectory given the input $x$ factorized as follows:
$p(\tau|x)=p\left(\{g_{t}\}_{t=1}^{T}|x\right)p(\hat{y}|x,\{g_{t}\}_{t=1}^{T})=\left(\prod_{t=1}^{T}p(g_{t}|x_{,}\{g_{l}\}_{l=1}^{t-1})\right)p(\hat{y}|x,\{g_{t}\}_{t=1}^{T}),$ (2)
Here, each intermediate step $g_{t}$ explicitly depends on the task description $x$ and all previously generated guidance steps, while the final prediction $\hat{y}$ is conditioned on the full set of guidance as well as the original input. Given the conditional trajectory distribution $p(\tau|x)$, we aim to maximize the likelihood gap between high-quality trajectories ($\tau^{+}$) and lower-quality trajectories ($\tau^{-}$). Formally this motivates the optimization of the contrastive objective:
$\max_{p(y,\{g_{t}\}_{t=1}^{T}|x),\zeta\geq 0}\mathbb{E}_{\tau^{+},\tau^{-}}\left[\log p(\tau_{+}|x)-\log p(\tau_{-}|x)-\zeta\right],$ (3)
where $\tau^{+}$ represents trajectories associated with desired outcomes, and $\tau^{-}$ denotes trajectories yielding suboptimal predictions. The margin term $\zeta\geq 0$ enforces a minimal desired gap between the two trajectory distributions, thus ensuring robust separation and effectively guiding the joint model toward generating improved intermediate guidance and predictions. Thus, we utilize the final correctness of the black-box LLM’s output $\hat{y}$ to evaluate the quality of the trajectory $u(\tau)$ as the reward of the intermediate guidance produced by the white-box LLM controller (Figure 2):
$u(\tau)=\text{eval}(\hat{y},y),$ (4)
[…]
@@
For initialization, a prompt $x$ is sampled from task space $\mathcal{X}$ and serves as the initial state $s_{0}=x$. At each subsequent step $t\in[T]$, the controller generates prompts $a_{t}$ based on the current $s_{t-1}$. In response to the controller’s action, the environment first initiates an inner-loop multi-turn interaction, where the black-box LLM iteratively adjusts execution steps based on the controller’s guidance and returns an observation $o_{t}$ based on the history $s_{t-1}$ and the current action $a_{t}$. If the problem remains unsolved, the controller initiates an outer-loop multi-turn interaction, refining instructions through feedback to more effectively guide the LLM. Consequently, the state transitions are updated to include the new action and observation:
@@
The framework formulates a Markov Decision Process (MDP), which offers the potential for solving tasks that require long-horizon generations, including long-term planning and multi-step reasoning. By obtaining feedback from $\text{eval}(\cdot)$, we can conduct multi-turn optimization over the white-box LLM controller on the intermediate generations. Additionally, the multi-turn interaction with the environment during the data sampling stage can help improve data quality. Although optimizing this guidance presents challenges due to the inaccessibility of the black-box LLM’s parameters that preclude backpropagation of gradients during training, the existing reinforcement learning techniques, e.g.schulman2017proximal ; rosset2024direct , can be used for policy optimization.
3 Matryoshka Pilot (M-Pilot)
In this section, we specialize the white-box LLM controller that generates intermediate guidance to assist in task understanding and problem-solving in Section 3.1 and discuss the data collection procedure by interacting with black-box LLM in Section 3.2, which will be used for M-Pilot training to align the outputs of the black-box LLM with preferences in Section 3.3.
@@ 3.1 Instantiation of White-Box LLM Controller
We instantiate the white-box LLM as a controller to generate additional guidance that assists the black-box LLM in understanding and solving a diverse range of problems. Given varying complexity and distinct characteristics of different tasks, the controller should be capable of generating guidance in various formats. Examples of reasoning, planning, and personalization tasks are in Figure 3.
Figure 3: Examples of intermediate guidance generated by M-Pilot for complex reasoning, planning, and personalization tasks.
Problem Decomposition for Reasoning. For reasoning tasks, generating a sequence of reasoning steps is essential to solve the problem effectively. Existing works (zhouleast, ) have observed that models often perform poorly on tasks that require solving problems more complex than the exemplars provided in the prompts. To enable the model to better reason and overcome the easy-to-hard generalization issue, one strategy is to decompose complex problems into a series of simpler sub-problems and solve them sequentially. Therefore, for reasoning tasks, the white-box LLM controller outputs decomposed sub-tasks to assist the subsequent black-box LLM generator in enhancing its reasoning capabilities.
High-Level Plan for Planning. For planning tasks, LLMs are required to generate a sequence of actions that constitute a plan to solve the given problems. A common strategy (sun2024adaplanner, ; zhao2024epo, ) is to apply hierarchical planning for complex solutions, where a high-level planner decomposes the task into sub-goals, and a low-level planner generates a sequence of admissible actions corresponding to each specific sub-goal. To enhance the black-box LLM’s planning capabilities, we leverage the white-box controller to generate high-level plans as guidance for simplification.
User History Summarization for Personalization. For personalization tasks, LLMs are required to tailor outputs to individual users. Existing work (richardson2023integrating, ) accomplishes this by concatenating the user’s input query with a profile summarizing the user’s preferences and behavior patterns. To enhance the black-box LLM’s personalization capabilities, we utilize the white-box LLM controller to generate summaries of user histories. This approach enables black-box LLMs to better understand users and generate tailored content accordingly.